2. QUESTION:

A fair coin is tossed, and a fair die is thrown. Write down sample spaces for

(a) the toss of the coin;

(b) the throw of the die;

(c) the combination of these experiments.

Let A be the event that a head is tossed, and B be the event that an odd number is thrown. Directly from the sample space, calculate P(A ∩ B) and P(A ∪ B).


1
Expert's answer
2022-06-21T12:27:53-0400

(a) "\{\text{Head},\text{Tail}\}"

(b) "\{1,2,3,4,5,6\}"

(c) "\{(1\cap\text{Head}), (1\cap\text{Tail}),\dots,(6\cap\text{Head}), (6\cap\text{Tail})\}"


Clearly "\mathbb{P}(A)=\frac{1}{2}=\mathbb{P}(B)". We can assume that the two events are independent, so "\mathbb{P}(A\cap B)=\mathbb{P}(A)\mathbb{P}(B)=\frac{1}{4}". Alternatively, we can examine the sample space above and deduce that three of the twelve equally likely events comprise "A\cap B". Also, "\mathbb{P}(A\cup B)=\mathbb{P}(A)+\mathbb{P}(B)-\mathbb{P}(A\cap B)=\frac{3}{4}", where this probability can also be determined by noticing from the sample space that nine of twelve equally likely events comprise "A\cup B".